OpenStudy (anonymous):

7x2 - 12x + 16 = 0 -12x 7x2 16

OpenStudy (anonymous):

What are the directions for this? Just so I know Im giving you the correct way of doing it.

OpenStudy (anonymous):

Select the term that describes the linear portion in this quadratic equation. @aussy123

OpenStudy (anonymous):

Okay. So basically all you have to do is move over 12x. How do you do that?

OpenStudy (anonymous):

add 12

OpenStudy (anonymous):

close! Subtract.

OpenStudy (anonymous):

but the 12 is a negative

OpenStudy (anonymous):

Oh sorry I just noticed.

OpenStudy (anonymous):

Is thats a new Problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

2a-4=\[\frac{ a-1 }{ a-4}\] Make sure you simplify one side first.

OpenStudy (anonymous):

Then you multiply one a-4 over on both sides. It cancels out on the right.

OpenStudy (anonymous):

2a-4(a-4)=a-1 2a-4a+16=a-1 Add or subtract Like Terms -2a+16=a-1 +1 +1 _________________________________ -2a+17=a +2a +2a _________________ 17= 3a Now divide and simplify if possible.